You are going to select first and second place prizes out of 5 entries. How many possible outcomes are there? (Hint: order matters here)
Answer #1

Answers : There are total 5 entries and we are going to select first and second price out of 5 entries.

Since here, order matters and also there is no replacement

We can solve this problem by using permutation.

Here, n=Total number of entries=5

r= Number of chosen entries=2

Permutation formula= n!/(n-r)!=5!/(5-2)!=5!/3!

=(5×4×3×2×1)/(3×2×1)

n P r=(5×4)=20

Therefore total number of possible outcomes are 20